The function h(t)=200-16t represents the height of a ball in t seconds.
How many seconds does it take the object to hut the ground?

Answer by stanbon(75887)

The function h(t)=200-16t represents the height of a ball in t seconds.
How many seconds does it take the object to hut the ground?
======
Height is zero when the ball hits the ground.
---
Solve::
200-16t^2 = 0
200 = 16t^2
t^2 = 12.5
time = 3.54 seconds
-----------------
cheers,
Stan H.
---------------